# Human-Activity AGV Quality Assessment: A Benchmark Dataset and an Objective Evaluation Metric

### v1

@@ 3.4.3 MOS Analysis for Text Prompts
3.4.4 Semantic Distortion Identification
We counted all body parts appearing in the $3,200$ AGVs of the Human-AGVQA dataset and the distortions observed for each body part, as shown in Figure 5.
In Figure 5(a), we observe that the face, arms, and torso account for the largest proportions, while the legs and feet are significantly less than three parts. Notably, a small portion of the videos lack any visible body parts, indicating that some T2V models struggle to comprehend the text prompts and generate the corresponding human figures. In Figure 5(b), it is evident that the torso exhibits the fewest distortions among the five body parts, whereas the face, arms, and legs show more pronounced distortions. This suggests that current T2V models face significant challenges in generating complete and realistic body parts. The performance analysis of T2V models for distortion identification of human bodies can be found in the supplemental material.
4 Proposed Model
As illustrated in Figure 6, the proposed AI-generated human activity video quality (GHVQ) methods consists of four modules: a spatial quality analyzer, an action quality analyzer, a text feature extractor, and a quality regressor, which are detailed as follows:
4.1 Spatial Quality Analyzer
@@
4.1.1 Human-focused Quality Feature Extraction
For frame ${\bm{y}}_{i}$, we first apply a human body-part segmentation model [36], to explicitly detect the body-part masks $M_{i}$, where $M_{i}\in\mathbb{R}^{C_{m}\times H\times W}$ represents five segmentation masks corresponding to the face, arms, torso, legs, feet. Then, we design a shallow hourglass network [60] to extract the high-resolution feature maps $F_{h,i}$ that preserve the resolution of human appearance while capturing rich low-level features. Here, $F_{h,i}\in\mathbb{R}^{C_{h}\times H\times W}$, with $C_{h}$ representing the number of channels of $F_{h,i}$.
@@ 4.1.1 Human-focused Quality Feature Extraction
Feature Refinement and Fusion. For $F_{{\rm inner},i}$ and $F_{{\rm cross},i}$, we apply a CNN followed by a Transformer encoder [23] to further refine the inner-body and cross-body features:
$F^{\prime}_{{\rm inner},i}={\rm Transformer}({\rm Conv}(F_{{\rm inner},i})),$ (4)
@@
where $\rm Transformer$ denotes the Transformer encoder, and the $\rm Conv$ represents a two-layer CNN. Finally, we concatenate them and then apply a CNN layer followed by a self-attention layer [66] to derive the human-focused quality features:
$F_{{\rm bq},i}$ $={\rm Attention}({\rm Conv}([F^{\prime}_{{\rm inner},i},F^{\prime}_{{\rm cross},i}])),$ (5)
@@
4.1.2 Holistic Quality Feature Extraction
To extract quality-aware features of the entire frame, we pre-train a ViT [15] model on Pick-a-Pic [37], a large-scale text-to-image (T2I) quality assessment dataset, to learn the quality-aware feature representation for AIGC images. Then, for frame ${\bm{y}}_{i}$, we employ the pre-trained ViT [15] as the feature extractor to compute the holistic quality features:
$F_{{\rm hq},i}={\rm{ViT}}({\bm{y}}_{i}).$ (6)
Table 1: The performance comparison of existing IQA-based, AQA-based, and VQA-based metrics with our proposed GHVQ metrics on the Human-AGVQA dataset. NSC refers to natural scene content.
@@ 4.1.2 Holistic Quality Feature Extraction
Human Appearance NIQE (ISPL, 2012) [58] NA (handcraft) NSC 0.234 0.273 0.564 0.583
BRISQUE (TIP, 2012) [57] NA (handcraft) NSC 0.261 0.314 0.593 0.629
HyperIQA (CVPR, 2020) [68] TID2013 [63] NSC 0.314 0.339 0.684 0.705
UNIQUE (TIP, 2021) [101] KonIQ-10K [32] NSC 0.243 0.280 0.671 0.695
MUSIQ (ICCV, 2021) [34] KonIQ-10K [32] NSC 0.229 0.262 0.584 0.609
StairIQA (JSTSP, 2023) [72] KonIQ-10K [32] NSC 0.334 0.393 0.665 0.671
CLIP-IQA (AAAI, 2023) [81] KonIQ-10K [32] NSC 0.329 0.368 0.672 0.683
LIQE (CVPR, 2023) [102] KonIQ-10k [32] NSC 0.285 0.332 0.641 0.664
MA-AGIQA (ACMMM, 2024) [84] AGIQA-3k [43] AIGC 0.377 0.380 0.728 0.735
Q-Align (ICML, 2024) [89] fused  [32][17][96][21] NSC 0.364 0.419
GHVQ (proposed) AIGC —- 0.812 0.821
Action Continuity ACTION-NET (ACMMM, 2020) [99] Kinetics [6] NSC 0.197 0.221 0.541 0.551
USDL (CVPR, 2020) [74] Kinetics [6] NSC 0.207 0.260 0.583 0.582
CoRe (ICCV, 2021) [97] Kinetics [6] NSC 0.179 0.213 0.560 0.561
TSA (CVPR, 2022) [92] Kinetics [6] NSC 0.202 0.258 0.607 0.611
Motion Smoothness (CVPR, 2024) [33] AMT [48] AIGC 0.249 0.30
Temporal Flickering (CVPR, 2024) [33] RAFT [75] AIGC 0.109 0.238
Action-Score (CVPR, 2024) [54] VideoMAE V2 [82] AIGC 0.208 0.284
Flow-Score (CVPR, 2024) [54] RAFT [75] AIGC 0.153 0.255
GHVQ (proposed) AIGC 0.788 0.792
Overall Video TLVQM (TIP, 2019) [38] NA (handcraft) NSC 0.275 0.314 0.613 0.624
RAPIQUE (JSP, 2021) [78] NA (handcraft) NSC 0.313 0.351 0.632 0.648
VIDEAL (TIP, 2021) [77] NA (handcraft) NSC 0.338 0.357 0.638 0.651
PatchVQ (CVPR, 2021) [96] LSVQ [96] NSC 0.378 0.404 0.667 0.703
SimpleVQA (ACMMM, 2022) [71] LSVQ [96] NSC 0.368 0.380 0.695 0.709
FastVQA (ECCV, 2023) [87] LSVQ [96] NSC 0.385 0.408 0.708 0.733
DOVER (ICCV, 2023) [88] DIVIDE [88] NSC 0.315 0.337 0.713 0.736
T2VQA (ACMMM, 2024) [39] T2VQA [39] AIGC 0.357 0.364 0.751 0.754
UGVQ (Arxiv, 2024) [103] LGVQ [103] AIGC 0.349 0.358 0.748 0.757
EvalCrafter (CVPR, 2024) [54] DIVIDE [88] NSC 0.328 0.341
Q-Align (ICML, 2024) [89] fused  [32][17][96][21] AIGC 0.427 0.485
GHVQ (proposed) AIGC 0.780 0.788
@@
4.2 Action Quality Analyzer
The action quality analyzer is designed to quantify temporal continuity of videos, serving as a complement to the spatial quality analyzer in assessing comprehensively the quality of AGVs. To address this, we utilize an action recognition model, as the action quality analyzer to capture the temporal continuity of human body movement. Since action recognition models are trained on large-scale, authentic-captured human action datasets, such as the Kinetics series [6], they can effectively extract motion representations of human bodies. Moreover, previous VQA studies [71] have also demonstrated that these features are useful for video quality evaluation. Specifically, for video $\bm{x}$ and the action recognition network $\rm{SlowFast}$ [19], we calculate the action quality features $F_{\rm aq}$ using all frames in $\bm{x}$.
4.3 Text Feature Extractor
Since the purpose of T2V models is to generate video content that aligns with the text prompt, it is necessary to measure the alignment between the video content and the text prompt. Following most T2I and T2V alignment evaluation studies [33, 84], we utilize the text encoder of CLIP [37] to extract semantic features $F_{\rm tq}$ of the text prompt $p$.
4.4 Quality Regressor
@@
The loss function consists of the mean absolute error (MAE) loss, rank loss [39], and binary cross-entropy loss. The MAE loss, and rank loss are used to optimize the quality scoring task, and the binary cross-entropy loss is used to optimize the distortion identification task.
5 Experiments
5.1 Experiment Settings
Compared Quality Metrics. Since no specific metrics have been designed to evaluate human appearance quality, we compared $10$ IQA models to assess the quality of video frames as a proxy for human appearance. For action continuity quality and overall video quality, we select $8$ action quality assessment (AQA) and $11$ VQA methods as comparison respectively. For body parts detection and their distortion identification, there are also no specific metrics, we benchmark $9$ general video large multi-modality (LMM) model for this task. A detailed introduction to these methods can be found in the supplemental material.
Dataset Splits. We split the Human-AGVQA dataset into $70$% for training, $10$% for validation, and $20$% for test. For video quality scoring, the compared metrics are evaluated on the test set using two approaches: zero-shot testing and fine-tuning testing. For body parts detection and distortion identification, we directly test zero-shot performance of compared LMMs.
@@ 5.1 Experiment Settings
5.2 Performance Comparison
5.2.1 Quality Scoring Task
@@
Human Appearance Quality. We observe that all compared IQA methods perform poorly in evaluating human appearance quality in the zero-shot setting, which highlights that distortions in human appearance in AGVs represent entirely new artifacts that existing metrics cannot handle. Fine-tuning on the Human-AGVQA dataset improves the performance of data-driven methods, but their results remain suboptimal. This may be because these methods focus on learning feature representations from global images, rather than human-focused regions. The proposed GHVQ achieves the best performance, outperforming the second-best method by $0.812$ and $0.821$ in terms of SRCC and PLCC, demonstrating the effectiveness of its design for extracting human appearance-focused features.
Table 2: The accuracy of video LMM models in identifying the occurrences and distortions of body parts in the Human-AGVQA dataset.
Models Occurrences / Distortions
Face Arms Torso Legs Feet Avg
LLaMAVID (7B) [47] 91.9 / 56.6 82.9 / 62.1 81.8 / 46.5 37.2 / 30.7 25.1 / 41.5 63.8 / 47.7
VideoChatGPT (7B) [56] 73.9 / 71.7 79.7 / 76.2 75.2 / 37.6 38.8 / 23.0 29.2 / 19.4 59.4 / 45.6
VideoLLaMA2 (7B) [11] 22.5 / 24.9 14.1 / 14.9 41.2 / 61.2 61.9 / 75.8 70.4 / 78.3 42.2 / 51.0
VILA1.5 (7B) [51] 22.4 / 36.6 20.1 / 22.5 16.5 / 60.7 60.8 / 74.7 71.3 / 76.8 38.2 / 54.3
NeXT-Video (7B) [44] 62.5 / 37.6 56.8 / 22.5 78.6 / 60.9 58.2 / 75.0 65.3 / 77.6 64.3 / 54.7
OneVison (7B) [42] 61.9 / 29.8 71.9 / 19.8 64.6 / 62.2 52.2 / 77.0 67.2 / 80.3 63.6 / 53.8
Qwen2-VL (7B) [83] 63.1 / 44.3 57.9 / 25.7 71.0 / 60.4 62.6 / 76.1 74.1 / 79.9 70.3 / 57.3
GPT-4o  [61] 92.4 / 29.7 89.7 / 19.3 85.1 / 61.6 62.1 / 77.4 76.1 / 80.1 80.0 / 53.6
GPT-4o-mini  [61] 90.6 / 28.4 86.5 / 20.1 82.7 / 60.4 62.6 / 75.3 74.0 / 77.3 78.6 / 52.3
GHVQ (proposed) 92.6 / 72.3 87.1 / 80.6 87.1 / 63.4 63.5 / 78.1 77.2 / 81.5 81.5 / 75.2
@@
Action Continuity Quality. Similar to human appearance quality, the compared AVA methods and temporal metrics of AIGC VQA methods show a very low correlation in assessing action continuity quality. This is because AVA methods are typically designed to quantify the completeness of professional sports performances, while the temporal metrics in AIGC VQA methods rely on handcrafted motion descriptors that do not capture semantic content, such as those involving individuals. In the fine-tuning results, GHVQ outperforms the re-trained AVA methods, surpassing the second-best method by $0.788$ and $0.792$ in terms of SRCC and PLCC, demonstrating that the proposed model architecture is better suited to measure action continuity quality.
@@
Overall Video Quality. The performance of VQA metrics in assessing overall video quality is slightly superior to that of other metrics for human appearance and action continuity quality. Generally, we find that AIGC-based VQA methods outperform general data-driven VQA methods, and both of these are more effective than knowledge-driven VQA methods. Similarly, re-training these methods leads to significant improvements. Since GHVQ comprehensively extracts human-focused, action continuity, and holistic frame quality features, the proposed GHVQ outperforms existing VQA methods, improving performance by $0.780$ and $0.788$ in terms of SRCC and PLCC.

### v2

@@ 3.4.3. MOS Analysis for Text Prompts
Figure 6. The GHVQ framework comprises four essential modules: a spatial quality analyzer for extracting human-focused and holistic quality features, an action quality analyzer for capturing temporal continuity of the body movement, a text feature extractor for identifying video content that needs to be generated by text prompts, and a quality regressor for mapping these quality-aware features to multi-dimensional quality scores and semantic distortion labels.
3.4.4. Semantic Distortion Identification
We counted all body parts appearing in the $6,000$ AGVs of the Human-AGVQA dataset and the distortions observed for each body part, as shown in Figure 5.
In Figure 5(a), we observe that the face, hands, arms, and torso account for the largest proportions, while the legs and feet are significantly less than four parts. Notably, a small portion of the videos lack any visible body parts, indicating that some T2V models struggle to comprehend the text prompts and generate the corresponding human figures. In Figure 5(b), it is evident that the torso exhibits the fewest distortions among the six body parts, whereas the hands, face and feet show more pronounced distortions. This suggests that current T2V models face significant challenges in generating complete and realistic body parts. The performance analysis of T2V models for distortion identification of human bodies can be found in the supplemental material.
4. Proposed Model
As illustrated in Figure 6, the proposed AI-generated human activity video quality (GHVQ) method consists of four modules: a spatial quality analyzer, an action quality analyzer, a text feature extractor, and a quality regressor, which is detailed as follows:
4.1. Spatial Quality Analyzer
@@
4.1.1. Human-focused Quality Feature Extraction
For frame ${\bm{y}}_{i}$, we first apply a human body-part segmentation model (Khirodkar et al., 2024), to explicitly detect the body-part masks $M_{i}$, where $M_{i}\in\mathbb{R}^{C_{m}\times H\times W}$ represents six segmentation masks corresponding to the face, arms, torso, legs, feet. Then, we design a shallow hourglass network (Newell et al., 2016) to extract the high-resolution feature maps $F_{h,i}$ that preserve the resolution of human appearance while capturing rich low-level features. Here, $F_{h,i}\in\mathbb{R}^{C_{h}\times H\times W}$, with $C_{h}$ representing the number of channels of $F_{h,i}$.
@@ 4.1.1. Human-focused Quality Feature Extraction
Feature Refinement and Fusion. For $F_{{\rm inner},i}$ and $F_{{\rm cross},i}$, we apply a CNN followed by a Transformer encoder (Han et al., 2021) to further refine the inner-body and cross-body features:
(4) $F^{\prime}_{{\rm inner},i}={\rm Transformer}({\rm Conv}(F_{{\rm inner},i})),$
@@
where $\rm Transformer$ denotes the Transformer encoder, and the $\rm Conv$ represents a two-layer CNN. Finally, we concatenate them and then apply a CNN layer followed by a self-attention layer (Shaw et al., 2018) to derive the human-focused quality features:
(5) $F_{{\rm bq},i}$ $={\rm Attention}({\rm Conv}([F^{\prime}_{{\rm inner},i},F^{\prime}_{{\rm cross},i}])),$
@@
4.1.2. Holistic Quality Feature Extraction
To extract quality-aware features of the entire frame, we pre-train a ViT (Dosovitskiy et al., 2021) model on Pick-a-Pic (Kirstain et al., 2023), a large-scale text-to-image (T2I) quality assessment dataset, to learn the quality-aware feature representation for AIGC images. Then, for frame ${\bm{y}}_{i}$, we employ the pre-trained ViT (Dosovitskiy et al., 2021) as the feature extractor to compute the holistic quality features:
(6) $F_{{\rm hq},i}={\rm{ViT}}({\bm{y}}_{i}).$
Table 1. The performance comparison of existing IQA-based, AQA-based, and VQA-based metrics with our proposed GHVQ metrics on the Human-AGVQA dataset. NSC refers to natural scene content.
@@ 4.1.2. Holistic Quality Feature Extraction
Human Appearance NIQE (ISPL, 2012) (Mittal et al., 2012b) NA (handcraft) NSC 0.233 0.273 0.564 0.570
BRISQUE (TIP, 2012) (Mittal et al., 2012a) NA (handcraft) NSC 0.258 0.314 0.594 0.619
CNNIQA (CVPR, 2014) (Kang et al., 2014) TID2013 (Ponomarenko et al., 2015) NSC 0.243 0.296 0.620 0.646
HyperIQA (CVPR, 2020) (Su et al., 2020) TID2013 (Ponomarenko et al., 2015) NSC 0.318 0.339 0.684 0.696
UNIQUE (TIP, 2021) (Zhang et al., 2021) KonIQ-10K (Hosu et al., 2020) NSC 0.239 0.280 0.668 0.684
MUSIQ (ICCV, 2021) (Ke et al., 2021) KonIQ-10K (Hosu et al., 2020) NSC 0.224 0.262 0.583 0.601
StairIQA (JSTSP, 2023) (Sun et al., 2023) KonIQ-10K (Hosu et al., 2020) NSC 0.339 0.393 0.661 0.663
CLIP-IQA (AAAI, 2023) (Wang et al., 2023a) KonIQ-10K (Hosu et al., 2020) NSC 0.327 0.368 0.672 0.674
LIQE (CVPR, 2023) (Zhang et al., 2023b) KonIQ-10k (Hosu et al., 2020) NSC 0.283 0.332 0.640 0.656
MA-AGIQA (ACMMM, 2024) (Wang et al., 2024b) AGIQA-3k (Li et al., 2023a) AIGC 0.373 0.380 0.724 0.733
Q-Align (ICML, 2024) (Wu et al., 2023c) fused  (Hosu et al., 2020)(Fang et al., 2020)(Ying et al., 2021)(Gu et al., 2018) NSC 0.362 0.419 0.725 0.731
GHVQ (proposed) AIGC —- 0.805 0.809
Action Continuity ACTION-NET (ACMMM, 2020) (Zeng et al., 2020) Kinetics (Carreira and Zisserman, 2017a) NSC 0.198 0.223 0.541 0.553
USDL (CVPR, 2020) (Tang et al., 2020) Kinetics (Carreira and Zisserman, 2017a) NSC 0.208 0.259 0.583 0.591
CoRe (ICCV, 2021) (Yu et al., 2021) Kinetics (Carreira and Zisserman, 2017a) NSC 0.177 0.210 0.562 0.577
TSA (CVPR, 2022) (Xu et al., 2022) Kinetics (Carreira and Zisserman, 2017a) NSC 0.204 0.256 0.602 0.613
Motion Smoothness (CVPR, 2024) (Huang et al., 2024) AMT (Li et al., 2023b) AIGC 0.250 0.275
Temporal Flickering (CVPR, 2024) (Huang et al., 2024) RAFT (Teed and Deng, 2020) AIGC 0.137 0.239
Action-Score (CVPR, 2024) (Liu et al., 2024a) VideoMAE V2 (Wang et al., 2023b) AIGC 0.209 0.244
Flow-Score (CVPR, 2024) (Liu et al., 2024a) RAFT (Teed and Deng, 2020) AIGC 0.254 0.279
GHVQ (proposed) AIGC 0.771 0.778
Overall Video TLVQM (TIP, 2019) (Korhonen, 2019) NA (handcraft) NSC 0.272 0.312 0.603 0.617
RAPIQUE (JSP, 2021) (Tu et al., 2021b) NA (handcraft) NSC 0.313 0.351 0.621 0.637
VIDEAL (TIP, 2021) (Tu et al., 2021a) NA (handcraft) NSC 0.342 0.353 0.628 0.642
PatchVQ (CVPR, 2021) (Ying et al., 2021) LSVQ (Ying et al., 2021) NSC 0.379 0.399 0.657 0.693
SimpleVQA (ACMMM, 2022) (Sun et al., 2022) LSVQ (Ying et al., 2021) NSC 0.364 0.378 0.687 0.698
BVQA (TCSVT, 2022) (Li et al., 2022b) KoNViD-1k (Hosu et al., 2017) NSC 0.349 0.374 0.672 0.702
FastVQA (ECCV, 2023) (Wu et al., 2022) LSVQ (Ying et al., 2021) NSC 0.390 0.412 0.698 0.711
DOVER (ICCV, 2023) (Wu et al., 2023b) DIVIDE (Wu et al., 2023b) NSC 0.312 0.337 0.705 0.717
T2VQA (ACMMM, 2024) (Kou et al., 2024) T2VQA (Kou et al., 2024) AIGC 0.359 0.367 0.737 0.742
UGVQ (Arxiv, 2024) (Zhang et al., 2024) LGVQ (Zhang et al., 2024) AIGC 0.349 0.358 0.734 0.743
EvalCrafter (CVPR, 2024) (Liu et al., 2024a) DIVIDE (Wu et al., 2023b) AIGC 0.328 0.336
Q-Align (ICML, 2024) (Wu et al., 2023c) fused  (Hosu et al., 2020)(Fang et al., 2020)(Ying et al., 2021)(Gu et al., 2018) AIGC 0.422 0.481 0.715 0.723
GHVQ (proposed) AIGC 0.768 0.773
@@
4.2. Action Quality Analyzer
The action quality analyzer is designed to quantify the temporal continuity of videos, serving as a complement to the spatial quality analyzer in assessing comprehensively the quality of AGVs. To address this, we utilize an action recognition model, as the action quality analyzer to capture the temporal continuity of human body movement. Since action recognition models are trained on large-scale, authentic-captured human action datasets, such as the Kinetics series (Carreira and Zisserman, 2017a), they can effectively extract motion representations of human bodies. Moreover, previous VQA studies (Sun et al., 2022) have also demonstrated that these features are useful for video quality evaluation. Specifically, for video $\bm{x}$ and the action recognition network $\rm{SlowFast}$ (Feichtenhofer et al., 2019), we calculate the action quality features $F_{\rm aq}$ using all frames in $\bm{x}$.
4.3. Text Feature Extractor
Since the purpose of T2V models is to generate video content that aligns with the text prompt, it is necessary to measure the alignment between the video content and the text prompt. Following most T2I and T2V alignment evaluation studies (Huang et al., 2024; Wang et al., 2024b), we utilize the text encoder of CLIP (Kirstain et al., 2023) to extract semantic features $F_{\rm tq}$ of the text prompt $p$.
4.4. Quality Regressor
@@
The loss function consists of the mean absolute error (MAE) loss, rank loss (Wen and Wang, 2021), and binary cross-entropy loss. The MAE loss and rank loss are used to optimize the quality scoring task, and the binary cross-entropy loss is used to optimize the distortion identification task.
5. Experiments
Table 2. The accuracy of MLMM models in identifying the occurrences and distortions of body parts in the Human-AGVQA dataset.
Models Occurrences / Distortions
Face Hands Arms Torso Legs Feet Avg
LLaMAVID (7B) (Li et al., 2025) 83.25 / 38.24 26.94 / 31.99 68.61 / 36.00 81.72 / 46.40 57.91 / 30.60 24.99 / 21.39 57.24 / 34.10
VideoChatGPT (7B) (Maaz et al., 2023) 67.11 / 61.58 20.09 / 51.13 51.75 / 49.65 64.52 / 37.54 41.36 / 22.93 29.11 / 19.26 45.66 / 40.35
VideoLLaMA2 (7B) (Cheng et al., 2024) 26.70 / 24.77 20.64 / 14.82 21.64 / 20.29 54.41 / 31.13 21.74 / 45.72 30.33 / 48.24 29.24 / 30.83
VILA1.5 (7B) (Lin et al., 2024) 35.65 / 26.52 23.46 / 22.39 29.79 / 27.38 57.82 / 30.63 27.52 / 41.63 31.19 / 46.66 34.24 / 32.54
NeXT-Video (7B) (Li et al., 2024b) 60.97 / 42.56 35.96 / 27.21 45.24 / 31.18 64.48 / 51.80 32.22 / 34.85 45.23 / 57.53 47.35 / 40.86
OneVison (7B) (Li et al., 2024a) 62.21 / 44.06 39.21 / 39.05 46.17 / 33.50 66.90 / 53.13 33.27 / 40.93 47.11 / 50.17 49.15 / 43.47
Qwen2-VL (7B) (Wang et al., 2024a) 72.90 / 52.92 47.88 / 31.59 59.08 / 36.83 71.03 / 55.43 47.12 / 65.97 54.01 / 64.84 58.67 / 51.26
Qwen2.5-VL (7B) (Bai et al., 2025) 87.66 / 68.38 65.15 / 38.79 72.13 / 47.18 81.81 / 63.31 59.54 / 71.04 69.00 / 72.77 72.55 / 60.25
DeepSeek-VL2 (4.1B) (Wu et al., 2024) 71.80 / 53.56 44.81 / 33.38 57.45 / 51.60 82.67 / 50.67 43.89 / 62.97 60.95 / 52.76 60.26 / 50.82
GPT-4o  (OpenAI, 2024a) 92.33 / 29.63 82.28 / 24.95 88.51 / 19.23 84.95 / 60.31 62.04 / 73.31 75.96 / 71.02 81.01 / 46.41
GPT-4o-mini  (OpenAI, 2024a) 90.50 / 28.29 80.35 / 22.60 86.41 / 20.03 83.55 / 61.49 62.52 / 72.24 73.94 / 69.16 79.55 / 45.64
GHVQ (proposed) 92.99 / 78.36 84.78 / 62.86 88.50 / 70.25 90.23 / 73.13 64.26 / 83.71 77.37 / 87.35 83.02 / 75.94
5.1. Experiment Settings
Compared Quality Metrics. Since no specific metrics have been designed to evaluate human appearance quality, we compared $11$ IQA models to assess the quality of video frames as a proxy for human appearance. For action continuity quality and overall video quality, we select $8$ action quality assessment (AQA) and $12$ VQA methods as comparison respectively. For body parts detection and their distortion identification, there are also no specific metrics, we benchmark $11$ general video multi-modal large multi-modality (MLMM) model for this task. A detailed introduction to these methods can be found in the supplemental material.
Dataset Splits. We split the Human-AGVQA dataset into $70$% for training, $10$% for validation, and $20$% for test. For video quality scoring, the compared metrics are evaluated on the test set using two approaches: zero-shot testing and fine-tuning testing. For body parts detection and distortion identification, we directly test zero-shot performance of compared MLMMs.
@@ 5.1. Experiment Settings
5.2. Performance Comparison
5.2.1. Quality Scoring Task
@@
Human Appearance Quality. We observe that all compared IQA methods perform poorly in evaluating human appearance quality in the zero-shot setting, which highlights that distortions in human appearance in AGVs represent entirely new artifacts that existing metrics cannot handle. Fine-tuning on the Human-AGVQA dataset improves the performance of data-driven methods, but their results remain suboptimal. This may be because these methods focus on learning feature representations from global images, rather than human-focused regions. The proposed GHVQ achieves the best performance, outperforming the second-best method by $0.805$ and $0.809$ in terms of SRCC and PLCC, demonstrating the effectiveness of its design for extracting human appearance-focused features.
@@
Action Continuity Quality. Similar to human appearance quality, the compared AVA methods and temporal metrics of AIGC VQA methods show a very low correlation in assessing action continuity quality. This is because AVA methods are typically designed to quantify the completeness of professional sports performances, while the temporal metrics in AIGC VQA methods rely on handcrafted motion descriptors that do not capture semantic content, such as those involving individuals. In the fine-tuning results, GHVQ outperforms the re-trained AVA methods, surpassing the second-best method by $0.771$ and $0.778$ in terms of SRCC and PLCC, demonstrating that the proposed model architecture is better suited to measure action continuity quality.
@@
Overall Video Quality. The performance of VQA metrics in assessing overall video quality is slightly superior to that of other metrics for human appearance and action continuity quality. Generally, we find that AIGC-based VQA methods outperform general data-driven VQA methods, and both of these are more effective than knowledge-driven VQA methods. Similarly, re-training these methods leads to significant improvements. Since GHVQ comprehensively extracts human-focused, action continuity, and holistic frame quality features, the proposed GHVQ outperforms existing VQA methods, improving performance by $0.768$ and $0.773$ in terms of SRCC and PLCC.
5.2.2. Body Presence and Distortion Identifying Task
Identifying the presence and distortions of human body parts is a novel task, and we benchmark the performance of MLMM models for this task in Table 2. The key prompts that we utilized for the evaluation are listed below. Note that replace [body part] with one of the following: ”face”, ”hands”, ”arms”, ”torso”, ”legs”, or ”feet”. More details are shown in the supplementary material.
For body part occurrences:
# Is there a person’s [body part] appearing in this video? Answer with ”Yes” or ”No”. #.
For body part distortions:
# Is there any incompleteness, unrealistic appearance, or discontinuous movements of the person’s [body part] in this video? Answer with ”Yes” or ”No”. #.
As for MLMMS, the GPT-4 series performs best in terms of detecting the presence of body parts, while Qwen2.5-VL achieves the highest accuracy in identifying whether body parts exhibit semantic distortions. We observe that that identifying distortions in arms is the most challenging, followed by facial distortions while detecting distortions in legs and feet is relatively easier. This aligns with the inherent complexity of structures like faces and arms (particularly hands). Overall, our GHVQ model outperforms in both presence detection and distortion identification tasks, improving accuracy by $2.01\%$ and $15.69\%$, respectively.